The reversible chemical reaction A + B C + D has
Answer #1

Part A

at eqlibrium

[A] = [B] = 2.00-x

[C] = [D] =x

4.6 = x^2/(2.00-x)^2

square root

2.14 = x/2-x

4.28 - 2.14x = x

x= 1.36 M

[ A ] = 2.00-1.36 = 0.64 M

Part B

the reversible chemical reaction

A + B ⇌ C + D

1...2...0...0

1-x...2-x...x...x

so k = 4.6 = x^2/(1-x)*(2-x)

x = D = 1.4 M